Good Luck. (Need to show all your work to get full credit; no work, no credit) 1.For the given sample data below: a Construct a frequency distribution using six classes for a sample data that includes. Class Limits (5 pts), Class Boundaries (6 pts), Frequency (3 pts), Relative frequency (3 pts), Cumulative Frequency (3 pts), Midpoint X.5 pts), f-x-(5 pts), f. (6 pts). b. For constructed grouped data find: mean (8 pts), mode for grouped data (2 pts), modal class (2 pts), the variance (10 pts), and the standard deviation (2 pts) c. Draw histogram and discuss the shape of a distribution (10 pts), frequency polygon (10 pts) ogive (10 pts), and pie chart for the following table (10 pt). NOTE: you MUST use 2.0 as a lower limit of the first class and do appropriate estimation for class width and in other computations. 4.6 3.9 2.8 6.6 4.2 3.7 3.7 5.9 3.2 2.2 3.2 4.1 21 4.2 50 46 54 24 63 29 5.3 40 4.7 3.6 3.3 6.9 45 2.5 5.4 5.7 3.84.1 7.96.2 3.0 33 5.0 5.4 3.4 4.4 4.0 365.0 4.1 48 72 64 30 3.5 53 7.7 3.9 26 5.6 33 55 43 3.9 63 3.1 3.1 3.0 4.8 4.1 5 Class Limits Class boundaries Tally Frequency Relative CumuativeClas Frequency regueny
media%2Ff8b%2Ff8bdbf6e-73b9-4103-a2ae-1f
Answer #1

1. From the given data

Class Interval Class Boundaries Relative Cumulative Mid
LL UL LL UL f Frequency Rel. Freq. values (xi) fixi fixi^2
2 2.9 1.95 2.95 7 0.111111 0.11111111 2.45 17.15 42.0175
3 3.9 2.95 3.95 19 0.301587 0.41269841 3.45 65.55 226.1475
4 4.9 3.95 4.95 16 0.253968 0.66666667 4.45 71.2 316.84
5 5.9 4.95 5.95 12 0.190476 0.85714286 5.45 65.4 356.43
6 6.9 5.95 6.95 6 0.095238 0.95238095 6.45 38.7 249.615
7 7.9 6.95 7.95 3 0.047619 1 7.45 22.35 166.5075
Total = 63 280.35 1357.558

b)

2JA = 280.35= 4.45 63 Variance (c) =ー fx2-x =-1357. 558-4452 = 1.74603175 63 SD 、/variance = 1.32137495

c)

Histogram 19 20 18 16 14 12 10 16 12 Series1 6 4 C.l.1.32137495

2JA = 280.35= 4.45 63 Variance (c") =ー fx2-x =-1357. 558-4452 = 1.74603175 63 SD 、/variance = 1.32137495 Histogram 19 20 18 16 14 12 10 16 12 Series1 6 4 C.l.1.32137495